The largest constant magnetic field achieved in the laboratory is about 35.8 T. What is the magnetic energy density due to this field?
Answer #1

here,

magnetic field , B = 35.8 T

magnetic energy density = 0.50 * B^2/(u0)

magnetic energy density = 0.50 * 35.8^2/(4pi *10^-7)

magnetic energy density = 5.1 *10^8 J/m^3

the magnetic energy density is 5.1 *108 J/m^3